OpenStudy (coolme):

what does 6! mean in math?

OpenStudy (anonymous):

Six factorial. The product of all whole numbers from 1 thru 6.

OpenStudy (anonymous):

6*5*4*3*2*1*0

OpenStudy (anonymous):

?

OpenStudy (shadowfiend):

Not including 0.

OpenStudy (anonymous):

This is a serious website. people have questions that need to be answerd. If im mis understanding im sorry.

OpenStudy (shadowfiend):

I believe this is a legitimate question, vvnnhh.

OpenStudy (anonymous):

oh.. ahahaha

OpenStudy (anonymous):

6 factorial n! = n*(n-1)! 0! = 1 so f.e: 5! = 5*4! = 5 * 4 * 3! = 5 * 4 * 3 * 2! = 5 * 4 * 3 * 2 * 1! = 5 * 4 * 3 * 2 * 1 * 0! = 5 * 4 * 3 * 2 * 1 * 1

OpenStudy (anonymous):

Did he seriously ask what 6 means? i mean..

OpenStudy (anonymous):

does he mean 1 2 3 4 5 6?

OpenStudy (coolme):

is there are short way to find the answer? like 26!

OpenStudy (shadowfiend):

You missed an exclamation point :) 6 !, or 6 factorial.

OpenStudy (shadowfiend):

coolme: a calculator's about the only short way to figure that one out :/ Well, a calculator, Google, that kind of thing.

myininaya (myininaya):

i know a short cut for multiplying binomials lol

OpenStudy (coolme):

what is that??

myininaya (myininaya):

look up the pascal's triangle

myininaya (myininaya):

1 1 1 n=2 1 2 1 n=3 1 3 3 1 n=4 1 4 6 4 1 for (x+y)^n, the above are the coefficients (x+y)^2=x^2+2xy+y^2 (make sure each terms has exponents add up to be 2) (x+y)^3=x^3+3x^2y+3xy^2+y^3 (each term has exponents add up to be 3) (x+y)^4=x^4+4x^3y+6x^2y^2+4xy^3+y^4 ( each term as exponents add up be 4)

myininaya (myininaya):

we can continue the triangle for n=5 and so on... n=5: coefficients are 1 5 10 10 5 1 (x+y)^5=x^5+5x^4y+10x^3y^2+10x^2y^3+5xy^4+y^5